# Gaussian Differentially Private Human Faces Under a Face Radial Curve Representation

### v1

@@ A.3 Reparameterization
To optimally register two surfaces we use the methods as described in Jermyn et al. [2017] which we summarize next. Suppose we have two surfaces $f_{1},f_{2}\in\mathcal{F}$ which we wish to optimally register over the parameterization group $\Gamma$. Here $\mathcal{F}$ is as in 2.2 $\mathcal{F}\ni f:D\rightarrow\mathbb{R}^{3}$ where $D$ is the unit disk. The action of $\Gamma$ on a surface is right composition, $f\circ\gamma$ for $\gamma\in\Gamma$, and reparameterizes the surfaces but does not change its image which in our case is the face. The authors in Jermyn et al. [2017] leverage a transformation of $f$ referred to as the square-root normal field (SRNF) defined as $q=\frac{n}{|n|^{1/2}}$ where $n$ is the normal vector $n=\frac{\partial f}{\partial u}\times\frac{\partial f}{\partial v}$. The corresponding action of $\Gamma$ on $q$ is then $(q,\gamma):=\sqrt{J(\gamma)}q\circ\gamma$. This transformation is necessary for an isometric action of $\Gamma$ but the details of this are lengthy and not necessary for our application.
@@
The authors define an energy function $E_{reg}:\Gamma\rightarrow\mathbb{R}_{\geq 0}$ to implement an iterative gradient descent method for the registration. The energy is defined as $E_{reg}(\gamma)=\|q_{1}-(\tilde{q}_{2},\gamma)\|^{2}$ where $\tilde{q}_{2}$ is the “current" stage of the surface being reparameterized. Here $\gamma$ is the incremental reparameterization with $\tilde{q}_{2}=({q}_{2},\gamma)$. Here $q_{1}$ would be our template surface from 2.2 and we register all surfaces to this template.
@@
Since the reparameterization is done in an iterative manner it results in incrementally improved registration. This method is iterative as the gradient is taken about the identity of $\Gamma$, $\gamma_{identity}$ and thus lives in the tangent space of this element. Suppose we have an orthonormal basis =$\mathcal{B}=\{b\}$ for the tangent space of the $\Gamma$ at the identity, $T_{id}(\Gamma)$ where each $b$ be a unit vector in the tangent space $T_{id}(\Gamma)$. The full gradient is given by $\sum_{b_{i}\in\mathcal{B}}\langle q_{1}-\tilde{q}_{2},d(b_{i},\gamma)\rangle_{2}b_{i}$ and we use the implementation mentioned in 2.2 and refer the interested reader to the cited materials.
@@
This reparameterization is costly due to both the dimension, iterative construction, and reliance on a basis. To alleviate some expense, we set the center of the surface as the tip of the nose as noted in A.6 with the triangulated mesh before we impose the disk parameterization. The reparameterization iterative method thus has to search over a smaller space as the center of the disk is pre-registered.
A.4 Data
@@
For a full description of how the data is collected we refer to Sero et al. [2019] but we summarize and emphasize the relevant points. All participants have the same neutral face expressions, so the data is not heterogeneous in terms of facial expression. Each face is captured with 7150 points which the authors refer to as quasi-landmarks. Further, these landmarks are registered across all individuals which Sero et al. [2019] refer to as “homologous." This registration is a point-wise correspondence across faces while the registration in A.3 is an entire disk correspondence. Similar to our approach the faces the faces are aligned using Procrustes Analysis, which finds the optimal rotation but againt considering the faces as a set of points not a disk-like surface.
@@ A.5 Supplemental Notes on DP
Both $\epsilon$ and $\delta$ are pre-specified parameters referred to as the privacy budget. When $\delta=0$, this is referred to as pure differential privacy. Differential privacy is an attribute of the random mechanism and roughly states that the distributions over adjacent datasets are not too different. To achieve approximate DP for functional data Mirshani et al. [2017] establish the following mechanism.
Theorem A.2 (Mirshani et al. [2017]).
@@
A.6 Möbius Transformation
The reparameterization defined in 2.1 is computationally expensive. To make the repameterization less expensive, we leverage the Möbius transformation of conformal maps. We give a high level idea of this transformation with more details available in Choi and Lui [2015, 2018], Choi et al. [2020] and implementation available at GitHub repository Choi [2020].
@@ A.7 Proof details
$P(PL\geq\epsilon)$ $=P\left(\log\left[\frac{\exp-\frac{1}{2\sigma}(\|h(D)\|_{\mathcal{H}}^{2}-2T_{D}(x))\}}{\exp-\frac{1}{2\sigma}(\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2T_{D^{\prime}}(x))\}}\right]\geq\epsilon\right)$
$=P\left(\log\left[\exp\{-\frac{1}{2\sigma}(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x)\}\right]\geq\epsilon\right)$
$=P\left(-\frac{1}{2\sigma^{2}}(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x)\geq\epsilon\right)$
@@
$P(-\frac{1}{2\sigma^{2}}(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x)\geq\epsilon)$
$=P(-\frac{1}{2\sigma^{2}}(-\|h(D)-h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x-h(D))\geq\epsilon)$
$\leq P(-\frac{1}{2\sigma^{2}}(-\Delta^{2}-2(T_{D}-T_{D^{\prime}})(x-h(D))\geq\epsilon)$
$=P((T_{D}-T_{D^{\prime}})(x-h(D))\geq\sigma^{2}\epsilon-\frac{\Delta^{2}}{2})$
$=P(\sigma\Delta Z\geq\sigma^{2}\epsilon-\frac{\Delta^{2}}{2})$
$=P(Z\geq\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma})=\Phi(-\frac{\sigma\epsilon}{\Delta}+\frac{\Delta}{2\sigma})$
@@
$P(PL\leq\epsilon)$ $=P\left(\log\left[\frac{\exp-\frac{1}{2\sigma}(\|h(D)\|_{\mathcal{H}}^{2}-2T_{D}(x))\}}{\exp-\frac{1}{2\sigma}(\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2T_{D^{\prime}}(x))\}}\right]\leq\epsilon\right)$
$=P(Z\leq\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma})=\Phi(\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma}).$
@@
A.8 Additional Results
In 5.1 we describe how to privatize a point-wise average face. Figure 8 displays a comparison of two private faces with $\mu_{T}=2$ and $\mu_{3}$, each from two angles. The two faces look fairly similar, but this is an artifact of the noise making the face rough and fuzzy. The face on the right column has a smoother outline as compared to the face on the left. In Table 1 we see that the MSE of the face on the right is nearly half of that of the face on the left.

### v2

@@ A.3 Reparameterization
To optimally register two surfaces we use the methods as described in Jermyn et al. (2017) which we summarize next. Suppose we have two surfaces $f_{1},f_{2}\in\mathcal{F}$ which we wish to optimally register over the parameterization group $\Gamma$. Here $\mathcal{F}$ is as in 2.2 $\mathcal{F}\ni f:D\rightarrow\mathbb{R}^{3}$ where $D$ is the unit disk. The action of $\Gamma$ on a surface is right composition, $f\circ\gamma$ for $\gamma\in\Gamma$, and reparameterizes the surfaces but does not change its image which in our case is the face. The authors in Jermyn et al. (2017) leverage a transformation of $f$ referred to as the square-root normal field (SRNF) defined as $q=\frac{n}{|n|^{1/2}}$ where $n$ is the normal vector $n=\frac{\partial f}{\partial u}\times\frac{\partial f}{\partial v}$. The corresponding action of $\Gamma$ on $q$ is then $(q,\gamma):=\sqrt{J(\gamma)}q\circ\gamma$. This transformation is necessary for an isometric action of $\Gamma$ but the details of this are lengthy and not necessary for our application.
@@
The authors define an energy function $E_{reg}:\Gamma\rightarrow\mathbb{R}_{\geq 0}$ to implement an iterative gradient descent method for the registration. The energy is defined as $E_{reg}(\gamma)=\|q_{1}-(\tilde{q}_{2},\gamma)\|^{2}$ where $\tilde{q}_{2}$ is the “current” stage of the surface being reparameterized. Here $\gamma$ is the incremental reparameterization with $\tilde{q}_{2}=({q}_{2},\gamma)$. Here $q_{1}$ would be our template surface from 2.2 and we register all surfaces to this template.
@@
Since the reparameterization is done in an iterative manner it results in incrementally improved registration. This method is iterative as the gradient is taken about the identity of $\Gamma$, $\gamma_{identity}$ and thus lives in the tangent space of this element. Suppose we have an orthonormal basis =$\mathcal{B}=\{b\}$ for the tangent space of the $\Gamma$ at the identity, $T_{id}(\Gamma)$ where each $b$ is a unit “vector” in the tangent space $T_{id}(\Gamma)$. The full gradient is given by $\sum_{b_{i}\in\mathcal{B}}\langle q_{1}-\tilde{q}_{2},d(b_{i},\gamma)\rangle_{2}b_{i}$ and we use the implementation mentioned in 2.2 and refer the interested reader to the cited materials.
@@
This reparameterization is costly due to the dimension, iterative construction, and reliance on a basis. To alleviate some expense, we set the center of the surface as the tip of the nose as noted in A.7 with the triangulated mesh before we impose the disk parameterization. The reparameterization iterative method thus has to search over a smaller space as the center of the disk is pre-registered.
A.4 Data Details
@@
For a full description of how the data is collected we refer to Sero et al. (2019) but we summarize and emphasize the relevant points. All participants have the same neutral face expressions, so the data is not heterogeneous in terms of facial expression. Each face is captured with 7150 points which the authors refer to as quasi-landmarks. Further, these landmarks are registered across all individuals which Sero et al. (2019) refer to as “homologous.” This registration is a point-wise correspondence across faces while the registration in A.3 is an entire disk correspondence. Similar to our approach the faces are aligned using Procrustes Analysis, which finds the optimal rotation but again considering the faces as a set of points not a disk-like surface.
@@ A.5 Supplemental Notes on DP
Both $\epsilon$ and $\delta$ are pre-specified parameters referred to as the privacy budget. When $\delta=0$, this is referred to as pure differential privacy. Differential privacy is an attribute of the random mechanism and roughly states that the distributions over adjacent datasets are not too different. To achieve approximate DP for functional data Mirshani et al. (2017) establish the following mechanism.
Theorem A.2 (Mirshani et al. (2017)).
@@
A.6 Private Point-wise Face
Here we have more details on the benchmark method. We have $n=1000$ faces each with 7150 points. Let $D=\{X_{i}\}$ and $X_{i}\in\mathbb{R}^{7150\times 3}$ be the set of faces; further let $X_{i}[k,l]$ be the $lth$ coordinate, ($x,y,z$), of the $kth$ point of the $ith$ face, $l\in\{1,2,3\}$, $k\in\{1,\dots,7150\}$, and $i\in\{1,\dots,1000\}$. The points are “registered” across all the faces i.e. $X_{i}[k,\cdot]$ and $X_{j}[k,\cdot]$ represent the same facial feature e.g. the tip of the nose.
We compute the point-wise mean of $D$, the faces, as $\bar{X}=1/n\sum_{i}X_{i}$ and display this in the right panel of Figure 4. We have a total budget of $\mu_{T}$ for the entire face which we need to split across all points, $\bar{X}[k,\cdot]$. We sanitize, independently, the 7150 points, $\bar{X}[k,\cdot]$, each of which has 3 coordinates $\bar{X}[k,l]$. To split the budget evenly, each point’s coordinate is allocated $\mu_{p}:=\sqrt{\mu_{T}^{2}/(7150\cdot 3)}$ of the budget. That is, we require $(7150\cdot 3)$ many mechanisms and their budget composition is $\sqrt{\sum\mu_{p}^{2}}$, since we are using $\mu$-GDP. For each coordinate of each point of the average, $\bar{X}[k,l]$, we compute sensitivity as $\Delta_{k,l}=\max_{i,j}|X_{i}[k,l]-X_{j}[k,l]|$. We add noise to each coordinate of each point as $\bar{X}[k,l]+\Delta_{k,l}/\mu_{p}\cdot z$ where $z\sim\mathcal{N}(0,1)$.
A.7 Möbius Transformation
The reparameterization defined in 2.1 is computationally expensive. To make the repameterization less expensive, we leverage the Möbius transformation of conformal maps. We give a high level idea of this transformation with more details available in Choi & Lui (2015; 2018); Choi et al. (2020) and implementation available at GitHub repository Choi (2020).
@@ A.8 Proof Details
$P(PL\geq\epsilon)$ $=P\left(\log\left[\frac{\exp\{-\frac{1}{2\sigma}(\|h(D)\|_{\mathcal{H}}^{2}-2T_{D}(x))\}}{\exp\{-\frac{1}{2\sigma}(\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2T_{D^{\prime}}(x))\}}\right]\geq\epsilon\right)$
$=P\left(\log\left[\exp\{-\frac{1}{2\sigma}\left(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x)\right)\}\right]\geq\epsilon\right)$
$=P\left(-\frac{1}{2\sigma^{2}}(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x))\geq\epsilon\right)$
@@
$P\left(-\frac{1}{2\sigma^{2}}(\|h(D)\|_{\mathcal{H}}^{2}-\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x))\geq\epsilon\right)$
$=P\left(-\frac{1}{2\sigma^{2}}\left(-\|h(D)-h(D^{\prime})\|_{\mathcal{H}}^{2}-2(T_{D}-T_{D^{\prime}})(x-h(D)\right)\geq\epsilon\right)$
$\leq P\left(-\frac{1}{2\sigma^{2}}\left(-\Delta^{2}-2(T_{D}-T_{D^{\prime}})(x-h(D))\right)\geq\epsilon\right)$
$=P\left((T_{D}-T_{D^{\prime}})(x-h(D))\geq\sigma^{2}\epsilon-\frac{\Delta^{2}}{2}\right)$
$=P\left(\sigma\Delta Z\geq\sigma^{2}\epsilon-\frac{\Delta^{2}}{2}\right)$
$=P(Z\geq\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma})=\Phi\left(-\frac{\sigma\epsilon}{\Delta}+\frac{\Delta}{2\sigma}\right)$
@@
$P(PL\leq\epsilon)$ $=P\left(\log\left[\frac{\exp\{-\frac{1}{2\sigma}(\|h(D)\|_{\mathcal{H}}^{2}-2T_{D}(x))\}}{\exp\{-\frac{1}{2\sigma}(\|h(D^{\prime})\|_{\mathcal{H}}^{2}-2T_{D^{\prime}}(x))\}}\right]\leq\epsilon\right)$
$=P\left(Z\leq\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma}\right)=\Phi\left(\frac{\sigma\epsilon}{\Delta}-\frac{\Delta}{2\sigma}\right).$
@@
A.9 Additional Results
In 5.1 we describe how to privatize a point-wise average face. Figure 8 displays a comparison of two private faces with $\mu_{T}=2$ and $\mu_{T}=3$, each from two angles. The two faces look fairly similar, but this is an artifact of the noise making the face rough and fuzzy. The face on the right column has a smoother outline as compared to the face on the left. In Table 1 we see that the MSE of the face on the right is nearly half of that of the face on the left.
@@ A.9 Additional Results
In the right panel of Figure 3 we show the coordinate curves of one face radial curve for one face; In Figure 9 we show the coordinate curves of one face radial curve for one hundred faces. We see in Figure 9 that the curves (and hence the features) are aligned across individuals; this empirically shows the alignment and registration achieved by the methodology in Section 2.2.
Refer to caption Refer to caption Refer to caption
Figure 9: The $x$, $y$, and $z$ coordinate curves, from left to right, of one face radial curve for one hundred individuals.
Figure 10: A pipeline diagram of our methodology. We generate disk parameterized surfaces which we register and align to a template face. We extract the face radial curves from the disk parametrization and produce a sanitized mean via our GDP functional data mechanism.